OpenStudy (mayankdevnani):

A 1000 kg vehicle moving with a speed of a 20 m/s is brought to rest in a distance of 50 metre by applying brakes : (i) Find the acceleration. (ii) Calculate the unbalanced force acting on the vehicle.

OpenStudy (mayankdevnani):

@ash2326 @UnkleRhaukus @experimentX @hartnn

OpenStudy (mayankdevnani):

@.Sam. @sauravshakya

OpenStudy (mayankdevnani):

@pratu043 @ghazi

OpenStudy (mayankdevnani):

plz.........help me

OpenStudy (unklerhaukus):

so the acceleration is in opposite direction to movement

OpenStudy (ghazi):

it is deceleration

OpenStudy (mayankdevnani):

ok!!!! then

hartnn (hartnn):

v=u+at=0, find t from here put it in s=ut+0.5at^2 find a

OpenStudy (mayankdevnani):

I think acceleration=\[(\frac{v-u}{t})\]

OpenStudy (mayankdevnani):

@hartnn but, i don't know acc.

OpenStudy (ghazi):

F=m*a v^2=2*a*s

OpenStudy (ghazi):

here V=20

hartnn (hartnn):

t=-20/a put this in s=ut+0.5at^2

hartnn (hartnn):

s=50,u=20

OpenStudy (ghazi):

\[(20)^2= 2*a*50\]

OpenStudy (ghazi):

then F= m*a

hartnn (hartnn):

@ghazi u getting a=4, but a must be negative as it is deceleration..

OpenStudy (ghazi):

@hartnn @experimentX i have considered V as initial velocity ....it's all about consideration :)

OpenStudy (mayankdevnani):

@mathslover

OpenStudy (mayankdevnani):

ok!! then, @experimentX

OpenStudy (ghazi):

@mayankdevnani i guess just calculation is left

OpenStudy (mayankdevnani):

ok!! @ghazi

OpenStudy (mayankdevnani):

but, what's about a @experimentX

OpenStudy (ghazi):

deceleration = \[4m/s^2\]

OpenStudy (mayankdevnani):

right... @ghazi

OpenStudy (mayankdevnani):

but how you can find it... @ghazi please tell me step by step....

OpenStudy (ghazi):

then F= 1000*4 N it is called as unbalanced cuz...if forces would be balanced then either a=0 or net force = 0

OpenStudy (mayankdevnani):

yaa... f=4000N

OpenStudy (mayankdevnani):

but please tell me step by step

OpenStudy (ghazi):

\[(20)^2=2*a*50......a =\frac{ 400 }{ 100 }=4\]

OpenStudy (mayankdevnani):

but, F=m*a v^2=2*a*s how can you find it

OpenStudy (ghazi):

see from equation\[U^2=2*a*S\] we will have acceleration then use value of that acceleration in F=m*a to find F....since you've been provided with m=1000 and you have a=4 from above equation

OpenStudy (ghazi):

@mayankdevnani does that help?

OpenStudy (mayankdevnani):

yaaaaa..... @ghazi

OpenStudy (mayankdevnani):

but, how can you find s

OpenStudy (ghazi):

S=distance =50 m

mathslover (mathslover):

m = 1000 kg u = 20 v = 0 s = 50 v^2 - u^2 = 2a*50 -400 = 100 a a = -4 m/s^2 F = 1000 kg * - 4 m/s^2 = -4000 N

mathslover (mathslover):

that "-" MUST NOT BE NEGLECTED... you can say that force = 4000 N in opposite directionn

mathslover (mathslover):

*direction

OpenStudy (ghazi):

or retarding force

OpenStudy (mayankdevnani):

@mathslover s=50 ??? where you got 50... please tell me @ghazi and @mathslover

mathslover (mathslover):

@mayankdevnani it was given

mathslover (mathslover):

Please see the question again

OpenStudy (ghazi):

see the distance covered before vehicle comes to stop = 50 m = S it is given in your question

OpenStudy (mayankdevnani):

ya.... ok!!!! thanks a lot much thanks to @ghazi

OpenStudy (ghazi):

:) YW

mathslover (mathslover):

:( no problem ... sorry I was late but feel lucky to have answer by @ghass1978 :)

mathslover (mathslover):

\[\huge{\color{blue}{\mathbb{Good}}\space \color{red}{\mathsf{Work}} \space \color{green}{\textbf{Ghazi}}}\]

OpenStudy (ghazi):

Oh my KUDOS @mathslover sorry i am bad over latex lol

mathslover (mathslover):

lol no problem ghazi .. enjoy that sweet KUDOS .. :)

OpenStudy (ghazi):

:) thanks